OpenStudy (anonymous):

Find a ⋅ b. a = <5, 2>, b = <4, 5>

OpenStudy (tylerd):

multiplying vectors?

OpenStudy (anonymous):

Dot product is when you take corresponding components (so 5&4; 2&5) and multiply them together. Then add these products.

OpenStudy (anonymous):

essentially you're doing (5*4)+(2*5)

OpenStudy (bradely):

30

OpenStudy (anonymous):

20,10

OpenStudy (bradely):

add two =30

OpenStudy (anonymous):

oh so after i multiply them im suppose to add them

OpenStudy (anonymous):

\[a\cdot b = \sum a_ib_i\] In your case, a = <5, 2>, b = <4, 5>, so, \[a\cdot b = a_1b_1+a_2b_2 = 5(4) + 2(5) = ...\]

OpenStudy (anonymous):

Note: a = <\(a_1, a_2, ...\)> b = <\(b_1, b_2, ...\)>